OpenStudy (anonymous):

Fan and Medal Solve 3x2 + 2x + 7 = 0. Round solutions to the nearest hundredth. x ≈ -1.90 and x ≈ 1.23 x ≈ -1.23 and x ≈ 1.90 x ≈ -0.44 and x ≈ -3.56 No real solutions

OpenStudy (anonymous):

Im pretty sure its C /// can some one double check... really how?

OpenStudy (anonymous):

@PlanSlam

OpenStudy (anonymous):

no your right its the 3rd one i did a mistake sorry

OpenStudy (anonymous):

Can you show me how you did it so I can make sure... @PlanSlam

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

need more help?

OpenStudy (anonymous):

v3x^2+2x+7=0 It doesnt factor, so try the quad. Formula b^2- 4ac is the discriminant, and it is -80, so there are no real solutions. 2)3x^2-x-4=0 this factors (3x-4)(x+1)= 0